OpenStudy (anonymous):

multiply:

Parth (parthkohli):

0 * 0 = 0 again ;P

OpenStudy (anonymous):

\[\sqrt{2} (x- \sqrt{5})\]

Parth (parthkohli):

Use distributive. \(\Large \color{Black}{\Rightarrow (\sqrt{2} \times x) - (\sqrt{2} \times \sqrt{5}) }\) \(\Large \color{Black}{\Rightarrow x\sqrt2 - \sqrt{10} }\)

OpenStudy (anonymous):

2*2=4 same as 2+2=4

OpenStudy (anonymous):

ty parth